OpenStudy (anonymous):

x/x-1=1/x+1+3/x^-1 solve using an appropriate technique?????

OpenStudy (anonymous):

x/x-1=x-1+3/x^2-1 x=x+2/x+1 x(x+1)=x+2 x^2+x=x+2 x^2=2 X=+root(2),-root(2) as we know x^2-1=(x+1)(x-1)

OpenStudy (anonymous):

Redvelvet understood

OpenStudy (anonymous):

I am so sorry sai1234 i am at work and my boss walked in...so is the answer x^-1=(x+1)(x-1)??